### v3

@@ 4.5 Dynamically Modulating Policy Dominance
where $T_{min}$ and $T_{max}$ are the minimum and maximum allowable temperatures and $g(t)=\sum_{t^{\prime}=0}^{t}r^{\prime}(s_{t^{\prime}},a_{t^{\prime}})$ is the accumulated shifted human-induced reward. A plot of $T_{\psi}$ is shown in the Figure 2.
@@
How does shifting (Equation (5)) enable two-way switching? A high accumulated shifted human-induced reward ($>\eta$) indicates the agent has obeyed the intent-specific policy $\pi_{\psi}$ disproportionately more, which flags the need to mitigate the influence of $\pi_{\psi}$. Since the temperature in Equation (6) is monotonic with respect to accumulated reward $r^{\prime}$, a higher accumulated $r^{\prime}$ outputs a higher temperature, and this reduces the influence of the $\pi_{\psi}$ in the policy fusion step. Conversely, if the accumulated reward $r^{\prime}$ $<\eta$, the temperature is lowered, which strengthens $\pi_{\psi}$. Therefore, this setup enables a two-way dynamic switching between the policies. Based on the threshold $\eta$, the influence of $\pi_{\psi}$ is reduced when the accumulated $r^{\prime}$ is high, and strengthened when it is too low. Our overall algorithm for personalisation is summarised in Algorithm 1.
Table 1: Table showing different personalisation modes and corresponding human and task objectives.
Env Mode Human objective Task objective
@@
Avoidance Avoid a region
Mixed Prefer a region and avoid another
@@
Highway Preference Prefer a lane Move as fast as possible
Avoidance Avoid a lane without colliding
Mixed Prefer a lane and avoid another with other vehicles
@@ 5 Experiments
We demonstrate the effectiveness of our approach in our custom 2D Navigation environment, Highway [13], Pong [27] depicted in Figure 3. Specifically, we conduct personalisation experiments in three scenarios: Avoidance, Preference and Mixed  [^1] Code for our implementation can be found here. For Pong, we only consider the Preference scenario. To learn the intent-specific policy, we sampled 2000 trajectories from saved trajectory data. Now we describe the environments as follows:
@@
Table 2: Results in 2D Navigation. Undesired/Desired region is the average visit out of a total of 20 time steps. Values marked in blue indicate a top 15% task performance score relative to the best baseline, and those in red indicate a top 15% measure relative to the baseline that best adheres to human preferences. The up arrow $\uparrow$ indicates higher is better, and the down arrow $\downarrow$ indicates lower is better. The result averaged over 10 seeds with each seed containing 300 episodes with $T_{\phi}=0.4$, $T_{min}=1$, $T_{max}=10$, and $\eta=0$.
Mode Method Desired region $\uparrow$ Undesired region $\downarrow$ Score $\uparrow$
Preference DQN $0.085\pm 0.005$ $1.000\pm 0.000$
RUDDER $6.546\pm 0.342$ $0.006\pm 0.005$
MORL $9.575\pm 0.021$ $000.000\pm 0$ 0.000
Static $3.788\pm 0.452$ $0.602\pm 0.048$
Dynamic (Ours) $1.459\pm 0.140$ $1.000\pm 0.000$
Mixed DQN $0.050\pm 0.003$ $0.183\pm 0.006$ $1.000\pm 0.000$
RUDDER $2.790\pm 0.311$ $000.000\pm 0$ 0.000 $0.007\pm 0.007$
MORL $0.192\pm 0.022$ $0.119\pm 0.016$ $0.396\pm 0.035$
Static $1.116\pm 0.318$ $000.000\pm 0$ 0.000 $0.894\pm 0.035$
Dynamic (Ours) $0.268\pm 0.058$ $000.000\pm 0$ 0.000 $1.000\pm 0.000$
Avoidance DQN - $0.094\pm 0.006$ $1.000\pm 0.000$
RUDDER - $000.000\pm 0$ 0.000 $0.292\pm 0.055$
MORL - $0.026\pm 0.021$ $0.894\pm 0.099$
Static - $0.005\pm 0.005$ $1.000\pm 0.000$
Dynamic (Ours) - $0.006\pm 0.005$ $1.000\pm 0.000$
Table 3: Results for the Highway environment. Undesired/Desired lane columns indicate the average number of visits in 50 time steps. Values marked in blue indicate a top 15% task performance score relative to the best baseline, and those in red indicate a top 15% measure relative to the baseline that best adheres to human preferences. $\uparrow$ indicates higher values are better and $\downarrow$ indicates lower values are better for the quantity specified in the column. The hyper-parameters chosen are $T_{\phi}=0.6,T_{max}=5,T_{min}=0.3,\eta=0$.
Mode Method Desired lane $\uparrow$ Undesired lane $\downarrow$ Hits $\downarrow$ Score $\uparrow$
Avoidance DQN - $10.59\pm 0.96$ $0.11\pm 0.01$ $39.59\pm 0.46$
RUDDER - $00.00\pm 0$ 0.00 $0.73\pm 0.11$ $18.43\pm 2.54$
MORL - $5.80\pm 1.07$ $0.14\pm 0.23$ $37.44\pm 0.52$
Dynamic (Ours) - $0.19\pm 0.11$ $0.09\pm 0.02$ $38.80\pm 0.51$
@@ 5.1 Results
We first demonstrate the pitfalls of static fusion as described in Section 4.4 and compare it against dynamic fusion in 2D Navigation. As the performance of the static fusion baseline is sensitive to $T_{\psi}$, we tested a range of temperature values and found $T_{\psi}=T_{max}/2$ as the best-performing agent across all three modes. As observed in Table 2, such a static fusion agent performed well in the Avoidance mode, exhibiting behaviours that avoid the undesired region, while achieving high scores. However, as seen in the Preference and Mixed modes, the static fusion agent incurs severe drops in its score and instead exhibits a high number of visits to the desired region. This is indicative of the over-dominance of the intent-specific policy component in the personalised policy. We confirmed this by examining the agent’s behaviour, which was found to repeatedly visit the desired region, ignoring the primary navigation task. Although the performance of static fusion could possibly be improved in individual modes of personalisation, it would require extensive tuning of the temperature hyperparameter, which renders this approach practically infeasible.
Regarding the other baselines, as expected, DQN prioritised task performance (Score), while RUDDER exhibited over-adherence to human preferences at the cost of task performance. The MORL baseline which is designed to balance multiple objectives was found to be less effective compared to our proposed dynamic fusion method in terms of both task and human objectives. We attribute this under-performance to an inherent issue in MORL, where it struggles to determine which objective to prioritise, particularly in situations where conflicts in action selection arise. This issue becomes apparent in the mixed case of 2D Navigation, where MORL attempts to find a global solution satisfying both objectives, whereas our dynamic policy fusion method enables policy corrections locally, wherever adherence to human preferences is deemed necessary. Overall, the dynamic fusion agent strikes a balance between the primary task and intent-specific policy, assimilating both aspects into the behaviour, irrespective of the mode of personalisation.
These results establish fundamental limitations of static fusion and other baselines and highlight the need for dynamic fusion in order to satisfy both human and task objectives.
We further examine the properties of dynamic fusion in the Highway and Pong environments in Tables 3 and 4 respectively with the results averaged over 10 seeds each with 20 episodes. Here, DQN and RUDDER perform as expected, prioritising the task and human objectives respectively. In Highway, as seen in Table 3, DQN maximises the score and minimises hits (collisions), whereas RUDDER exhibits behaviours avoiding the undesired (first) lane and favouring the desired (third) lane. Similarly, as indicated in Table 4 in Pong, while DQN maximises the task score at the cost of seldom visiting the desired region, RUDDER causes the agent to remain in the preferred region, consistent with human preference (but incurring low scores). When compared to the dynamic policy fusion, the performance of the MORL baseline falls short of achieving both objectives. This is evident from the higher frequency of undesired lane visits in both the Avoidance and Mixed modes, and the lower frequency of desired lane visits in the Preference and Mixed modes. These results are detailed in Table 3. A similar trend is observed in Pong, where the MORL baseline not only failed to score but also spent less time in the preferred region compared to our method. These results indicate that dynamic policy fusion, while sometimes incurring a small reduction in the optimal score (achieved by DQN), enables the agent to adhere to the required preference (be it in the Avoiding, Preference or Mixed modes) while completing the primary task at hand. Additional experiments on different $\alpha$ in the Highway environment are provided in supplementary Section G.
@@
Table 4: Table showing the result in the Pong environment with $\eta=0$. Values marked in blue indicate a top 15% task performance score relative to the best baseline, and those in red indicate a top 15% measure relative to the baseline that best adheres to human preferences. % Desired region is the average fraction (in percentage) of an episode the agent stayed within the desired region shown in Figure 3 (b).
Method % Desired region $\uparrow$ Score $\uparrow$
@@
Table 5: Results correspond to the Highway environment with different values of $\eta$. $\uparrow$ indicates higher values are better and $\downarrow$ indicates lower values are better for the quantity specified in the column. The hyper-parameters chosen are $T_{\phi}=0.6,T_{max}=5,T_{min}=0.3$.
Mode $\eta$ Desired lane $\uparrow$ Undesired lane $\downarrow$ Hits $\downarrow$ Score $\uparrow$
Avoidance 0 - $0.28\pm 0.08$ $0.06\pm 0.01$ $39.83\pm 0.59$
1 - $0.35\pm 0.12$ $0.08\pm 0.02$ $38.96\pm 0.87$
2 - $0.19\pm 0.11$ $0.09\pm 0.02$ $38.80\pm 0.51$
Preference 0 $25.14\pm 1.44$ $0.05\pm 0.02$ $40.02\pm 0.61$
1 $24.90\pm 1.52$ $0.10\pm 0.03$ $38.84\pm 0.84$
2 $29.91\pm 1.38$ $0.06\pm 0.01$ $39.27\pm 0.59$
Mixed 0 $21.52\pm 2.00$ $0.85\pm 0.33$ $0.12\pm 0.01$ $38.50\pm 0.43$
1 $23.96\pm 1.67$ $0.56\pm 0.13$ $0.07\pm 0.01$ $39.56\pm 0.45$
2 $25.21\pm 2.55$ $0.27\pm 0.06$ $0.07\pm 0.04$ $39.15\pm 0.49$
Table 6: Table depicts the sensitivity analysis of $T_{max}$ in Highway environment mixed case. The hyper-parameters chosen are $T_{\phi}=0.6,T_{min}=0.3$ and $\eta=0$. $\uparrow$ indicates higher values are better and $\downarrow$ indicates lower values are better for the quantity specified in the column.
[…]
10 $19.21\pm 1.42$ $2.02\pm 0.43$ $0.06\pm 0.02$ $40.12\pm 0.59$
15 $17.60\pm 0.90$ $3.42\pm 0.43$ $0.09\pm 0.02$ $39.64\pm 0.45$
25 $9.53\pm 0.82$ $4.65\pm 0.52$ $0.10\pm 0.02$ $39.51\pm 0.53$
35 $14.36\pm 0.65$ $5.32\pm 0.58$ $0.10\pm 0.02$ $40.03\pm 0.59$
Effect of $\eta$: In Table 5, we examine the effect of $\eta$ on the performance of an agent under dynamic fusion. In theory, a higher $\eta$ would make it more likely for $T_{\psi}$ to remain low (refer to Figure 2), and thus the intent-specific policy becomes sharper. So while fusing the policies together, the obtained personalised policy becomes dominated by intent-specific policy. Our observations in Table 3 are in agreement with this behaviour, as in all preference modes, as $\eta$ increases, the desired lane visit increases and the undesired lane visit decreases, suggesting that personalised policy is dominated by the intent-specific policy component. In all of our experiments, we choose $\eta=0$. However, an ideal choice of $\eta$ may deviate from this value in other environments.
Effect of $T_{max}$: The parameter $T_{max}$ represents the maximum temperature in Equation (6). Theoretically, as $T_{max}$ is increased, the intent-specific policy should weaken. The observation in Table 6 supports this notion that when $T_{max}$ is increased, the desired lane visit decreases and the undesired lane visit increases suggesting a declining influence of intent-specific policy.
@@ Appendix A Simulating human feedback
Our method involves collecting human feedback for trajectories used for training the task-specific policy. This is done by simulating humans. We conducted experiments in three personalisation modes such as preference, Avoidance and Mixed. In each case, we counted the number of times the agent met the personalisation criteria within a trajectory and this is regarded as the ground truth score for a trajectory. For instance, In 2D navigation, if the human wishes the agent to visit a preferred region, we count the number of times the agent visited that state and a positive score is given to that trajectory. Similarly, in the avoidance case, a negative score is given and in the mixed case, we sum both the positives and the negatives.
Applying this method in a real-world context could involve several challenges. Using a real human for scoring trajectories, for example, may have limitations with regards to the number of trajectories to be labeled. Fewer labeled trajectories may have an adverse effect on the performance of our approach. In addition to this, a real subject would be prone to fatigue and may provide more arbitrary score labels. For instance, their scores may be biased based on the quality of trajectories presented to them immediately prior to scoring. These, along with a number of other challenges may need to be overcome before applying this approach to real-world scenarios.

### v4

@@ 4.5 Dynamically Modulating Policy Dominance
where $T_{min}$ and $T_{max}$ are the minimum and maximum allowable temperatures and $g(t)=\sum_{t^{\prime}=0}^{t}r^{\prime}(s_{t^{\prime}},a_{t^{\prime}})$ is the accumulated shifted human-induced reward. A plot of $T_{\psi}$ is shown in the Fig 2.
@@
How does shifting (Equation (5)) enable two-way switching? A high accumulated shifted human-induced reward ($\eta$) indicates the agent has obeyed the intent-specific policy $\pi_{\psi}$ disproportionately more, which flags the need to mitigate the influence of $\pi_{\psi}$. Since the temperature in Equation (6) is monotonic with respect to accumulated reward $r^{\prime}$, a higher accumulated $r^{\prime}$ outputs a higher temperature, and this reduces the influence of the $\pi_{\psi}$ in the policy fusion step. Conversely, if the accumulated reward $r^{\prime}$ $\eta$, the temperature is lowered, which strengthens $\pi_{\psi}$. Therefore, this setup enables a two-way dynamic switching between the policies. Based on the threshold $\eta$, the influence of $\pi_{\psi}$ is reduced when the accumulated $r^{\prime}$ is high, and strengthened when it is too low. Our overall algorithm for personalisation is summarised in Algorithm 1.
Table 1: Different personalisation modes and their corresponding human and task objectives.
Environment Mode Human objective Task objective
@@
Avoidance Avoid a region Move to the target
Mixed Prefer a region and avoid another Move to the target
@@
Highway Preference Prefer a lane Move as fast as possible without colliding
Avoidance Avoid a lane Move as fast as possible without colliding
Mixed Prefer a lane and avoid another Move as fast as possible without colliding
@@ 5 Experiments
We demonstrate the effectiveness of our approach in our custom 2D Navigation environment, Highway [12], Pong [23] depicted in Fig 3. Specifically, we conduct personalisation experiments in three scenarios: Avoidance, Preference and Mixed  [^1] Code for our implementation can be found here. For Pong, we only consider the Preference scenario. To learn the intent-specific policy, we sampled 2000 trajectories from saved trajectory data. Now we describe the environments as follows:
Table 2: Results in 2D Navigation. Undesired/Desired region is the average visit out of a total of 20 time steps. Values marked in blue indicate a top 15% task performance score relative to the best baseline, and those in red indicate a top 15% measure relative to the baseline that best adheres to human preferences. $\uparrow$ indicates higher is better and $\downarrow$ indicates lower is better. Results are averaged over 10 seeds, with each seed containing 300 episodes. The hyperparameters used are $T_{\phi}=0.4$, $T_{\min}=1$, $T_{\max}=10$, and $\eta=0$.
Mode Method
Desired region
visits $\uparrow$
Undesired region
visits $\downarrow$
Score
$\uparrow$
Preference DQN $0.085\pm 0.005$ $1.000\pm 0.000$
RUDDER $6.546\pm 0.342$ $0.006\pm 0.005$
MORL $9.575\pm 0.021$ $0.000\pm 0.000$
Static $3.788\pm 0.452$ $0.602\pm 0.048$
Dynamic (Ours) $1.459\pm 0.140$ $1.000\pm 0.000$
Mixed DQN $0.050\pm 0.003$ $0.183\pm 0.006$ $1.000\pm 0.000$
RUDDER $2.790\pm 0.311$ $000\pm 0.000 0.0$$007\pm 0.007$
MORL $0.192\pm 0.022$ $0.119\pm 0.016$ $0.396\pm 0.035$
Static $1.116\pm 0.318$ $0.000\pm 0.000$ $0.894\pm 0.035$
Dynamic (Ours) $268\pm 0.058 0.0$$000\pm 0.000$ $1.000\pm 0.000$
Avoidance DQN $0.094\pm 0.006$ $1.000\pm 0.000$
RUDDER $0.000\pm 0.000$ $0.292\pm 0.055$
MORL $0.026\pm 0.021$ $0.894\pm 0.099$
Static $0.005\pm 0.005$ $1.000\pm 0.000$
Dynamic (Ours) $0.006\pm 0.005$ $1.000\pm 0.000$
@@
Table 3: Results for the Highway environment. Undesired/Desired lane columns indicate the average number of visits in 50 time steps. Values marked in blue indicate a top 15% task performance score relative to the best baseline, and those in red indicate a top 15% measure relative to the baseline that best adheres to human preferences. $\uparrow$ indicates higher values are better and $\downarrow$ indicates lower values are better for the quantity specified in the column. The hyper-parameters chosen are $T_{\phi}=0.6,T_{\max}=5,T_{\min}=0.3,\eta=0$.
Mode Method
Desired lane
visits $\uparrow$
Undesired lane
visits $\downarrow$
Hits
$\downarrow$
Score
$\uparrow$
Avoidance DQN $59\pm 0.96 0.11±0.$$1\pm 0.01$ $39.59\pm 0.46$
RUDDER $0.00\pm 0.00$ $0.73\pm 0.11$ $18.43\pm 2.54$
MORL $5.80\pm 1.07$ $0.14\pm 0.23$ $37.44\pm 0.52$
Dynamic (Ours) $19\pm 0.11 0.0$$09\pm 0.02$ $38.80\pm 0.51$
@@ 5.1 Results
We begin by illustrating the limitations of static fusion (Section 4.4) and comparing it with our proposed dynamic fusion in the 2D Navigation task.
In this environment, the metric “desired region visits” quantifies adherence to user preference. Conceptually, the desired region (as specified by the human intent) may represent a checkpoint that the user expects the agent to acknowledge en route to the target. If the agent remains in this region indefinitely, it indicates over-dominance of the intent-specific policy, preventing progress toward the actual goal. An ideal personalised policy should therefore acknowledge the desired region but continue to complete the navigation task. As shown in Table 2, our dynamic fusion achieves exactly this balance: the agent briefly visits the desired region to respect the user’s preference but then proceeds reliably toward the target. This stands in contrast to static fusion and RUDDER, both of which over-emphasise user intent and consequently fail to complete the task.
The performance of static fusion is also highly sensitive to the temperature parameter $T_{\psi}$. We evaluated several values and found $T_{\psi}=T_{\max}/2$ to be the most effective across all three modes. As shown in Table 2, static fusion performs reasonably well in the Avoidance mode, successfully avoiding undesired regions while maintaining strong task scores. However, in the Preference and Mixed modes, it exhibits degradation: the agent becomes trapped in the desired region, resulting in poor task performance. This behaviour again reflects the over-dominance of the intent-specific policy under static fusion.
@@
Table 4: Results in the Pong environment with $\eta=0$. Values marked in blue indicate a top 15% task performance score relative to the best baseline, and those in red indicate a top 15% measure relative to the baseline that best adheres to human preferences. % Desired region is the average fraction of an episode (in percentage) that the agent stayed within the desired region shown in Fig 3(b).
Method
Desired region
% $\uparrow$
Score
$\uparrow$
@@
Table 5: Result when the user intent directly conflicts with the task objective in Highway in the preference case. The result is averaged over 5 seeds.
Method
Desired lane
% $\uparrow$
Speed violations
% $\downarrow$
Hits
$\downarrow$
Score
$\uparrow$
DQN $18.2\pm 4.24$ $15.72\pm 1.07$ $0.04\pm 0.02$ $38.84\pm 0.51$
Dynamic (Ours) $57.14\pm 8.52$ $00\pm 0.00 0.0$$02\pm 0.02$ $34.92\pm 0.70$
Table 6: Result with binary feedback in Highway in the preference case. The result is averaged over 5 seeds.
Method
Desired lane
% $\uparrow$
Hits
$\downarrow$
Score
$\uparrow$
DQN $21.8\pm 0.98$ $0.07\pm 0.02$ $40.25\pm 0.64$
Dynamic-Binary feedback (Ours) $60.78\pm 13.66$ $0.06\pm 0.02$ $38.28\pm 0.58$
Table 7: Result with binary feedback in Pong in the preference case. The result is averaged over 5 seeds.
Method
Desired region %
% $\uparrow$
Score
$\uparrow$
DQN $8.00\pm 0.63$ $0.57\pm 0.01$
Dynamic-Binary feedback (Ours) $42.54\pm 5.03$ $0.54\pm 0.05$
The behaviour of the other baselines is consistent with expectations. DQN focuses solely on maximising the task score, while RUDDER overfits to human preferences at the expense of task success. MORL, which attempts to reconcile both objectives, is less effective than our method. In particular, MORL struggles in scenarios such as Mixed 2D Navigation, where conflicting objectives arise. By seeking a global compromise, MORL fails to respond to local preference requirements, whereas our dynamic fusion adapts on the fly, applying corrections when human adherence is necessary. Overall, our method achieves a robust trade-off, incorporating both task and preference signals across all modes of personalisation.
These findings highlight the limitations of static fusion and MORL, and establish the necessity of dynamic policy fusion to satisfy both human and task objectives simultaneously.
In the Highway environment (Table 3), the expected behaviours of the baselines are again evident: DQN prioritises task performance, achieving high scores and minimal collisions, while RUDDER strongly favours human intent, spending the majority of time in the preferred lane but incurring more collisions and lower task scores. MORL attempts to balance both but falls short, as seen from its relatively high number of undesired lane visits in Avoidance and Mixed modes, and fewer desired lane visits in Preference and Mixed modes.
By contrast, our dynamic fusion method consistently balances task success with user alignment. In Avoidance, it minimises collisions and rarely enters the undesired lane while maintaining scores close to DQN. In Preference, it spends significantly more time in the desired lane than DQN or MORL, without compromising speed or safety. In Mixed, it adheres to both the lane preference and avoidance objectives simultaneously while sustaining strong task performance. These results highlight that dynamic fusion enables local corrections—adjusting the influence of user preferences adaptively—rather than committing to a fixed global trade-off as in MORL.
Turning to the Pong environment (Table 4), we evaluate the Preference mode, where the user favours keeping the paddle within a central region. Here, DQN achieves the highest task score but spends little time in the preferred region, while RUDDER almost exclusively remains in the preferred region at the cost of completely failing the task. MORL performs somewhat in between, but still fails to achieve meaningful task performance. Our dynamic fusion method strikes a balance: the agent spends a substantial proportion of time in the preferred region while still maintaining competitive scores, demonstrating effective alignment with user intent without sacrificing overall task viability.
Together, these results across 2D Navigation, Highway, and Pong consistently confirm that dynamic fusion outperforms static fusion, RUDDER, and MORL by adaptively balancing task objectives with user intent. While occasionally incurring a small drop in optimal task score relative to DQN, our approach achieves meaningful personalisation, ensuring that agents respect user preferences while still completing the primary task. Visualisation of the trajectories are available here.
We further evaluated the robustness of our approach under two challenging settings:
What if user intent directly conflicts with the task objective? In the Highway conflict scenario (Table 5), the user preference was to remain in the third lane while not exceeding a specified speed limit, which directly opposes the task objective of maximising speed without collisions. As expected, DQN prioritises speed and achieves the highest task score but only partially adheres to the lane preference and often violates the speed constraint. In contrast, our dynamic fusion method successfully maintains the lane preference and respects the speed limit, while still avoiding collisions. Although this leads to a lower overall task score (due to reduced speed), the outcome is consistent with our theoretical guarantees (Theorem 1): when the user and task objectives diverge within a bounded range, the resulting fused policy will also deviate from the optimal task policy in a controlled, bounded manner.
What if user feedback is binary? We investigated the use of binary feedback to reduce annotation effort and cognitive load for users. Instead of providing graded scores, trajectories were simply labelled as 1 if they contained the preferred states for more than 25% of the time, and 0 otherwise. As shown in Table  6 (Highway) and Table  7 (Pong), our dynamic fusion method remains effective even under this weaker supervision, producing behaviour that aligns with preferences while sustaining competitive task performance. This suggests that our approach is robust to coarse feedback and can remain practical in real-world scenarios where fine-grained labels are costly or noisy.
Together, these additional studies highlight the flexibility of dynamic fusion: it handles conflicting objectives with a controlled drop in performance and also performs reliably under binary human feedback.
@@ Appendix A Simulating human feedback
Our method involves collecting human feedback for trajectories used for training the task-specific policy. This is done by simulating humans. We conducted experiments in three personalisation modes: preference, avoidance, and mixed. In each case, we counted the number of times the agent met the personalisation criteria within a trajectory, and this count is regarded as the ground truth score for that trajectory. For instance, in 2D navigation, if the human wishes the agent to visit a preferred region, we count the number of times the agent visited that state and a positive score is given to that trajectory. Similarly, in the avoidance case, a negative score is given, and in the mixed case, we sum both the positive and negative counts.
Applying this method in a real-world context may introduce additional considerations. A key practical aspect is how trajectories are presented to a human for labelling while upholding the claim of no additional environment interactions. This can be addressed by allowing the user to passively view replays (i.e., recordings) of trajectories that were already collected during the task policy’s training. Since the agent does not actively generate new state-action transitions, this replay process does not constitute additional interaction.
In practice, the number of trajectories that can be labelled by a human may be limited, and labelling effort could be influenced by factors such as fatigue or subjective bias. While our experiments used a noise-free simulated human, future work could incorporate techniques to mitigate such issues, for example by employing semi-supervised learning, active learning, or noise-robust feedback aggregation methods. We view these extensions as promising directions to further enhance the practicality of our approach in real-world scenarios.